OpenStudy (anonymous):

solve this system using elimination x+y=30 x-y=36

OpenStudy (anonymous):

add the equations together: x+y+x-y = 2x = 66 So, x = 33 Substituting x, we have: 33+y = 30 --> y = -3

OpenStudy (anonymous):

\[x+y=30\] \[x-y=36\] add both of the equations together, both the positive and negative 'y' cancel out so the equation would be left as \[2x=66\]. That's when you divide both sides by 2 to get \[x=33\]. Now just plug 'x' into any of the two original equations, like so, \[33+y=30\]. Subtract 33 from both sides to get, \[y=-3\] tada!